# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 5.5 Real-World Applications
Self-play is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks (Lewis et al., 2017). Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) (Wang et al., 2024). Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse & Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
@@
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020). Moreover, EvoPlay (Wang et al., 2023) leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator (Jumper et al., 2021) to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football (Liu et al., 2022b; Haarnoja et al., 2024), quadruped competition (Xiong et al., 2024), and multi-drone volleyball (Xu et al., 2025), with significant efforts dedicated to Sim2Real transitions for real-world success (Gao et al., 2023).
6 Conclusion
@@
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play that agents interact with copies or past versions of themselves has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the game theory background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. We provide detailed explanations on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within various scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Also, future work can focus on how to integrate self-play with LLMs and achieve real-world applications. In conclusion, self-play is vital to modern RL research, offering key insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.
Appendix A Appendix
A.1 Abbreviations in Alphabetical Order
ABR Approximate Best Response
AI Artificial Intelligence
ARM Advantage-based Regret Minimization
ARMAC Advantage Regret-Matching Actor-Critic
BR Best Response
CCE Coarse Correlated Equilibrium
CE Correlated Equilibrium
CFR Counterfactual Regret Minimization
COP Combinatorial Optimization Problem
CSPL Curriculum Self-Play Learning
CVRP Capacitated Vehicle Routing Problem
DCFR Discounted Counterfactual Regret Minimization
DO Double Oracle
EGTA Empirical Game-Theoretic Analysis
FP Fictitious Play
FSP Fictitious Self-Play
FTW For The Win
FXP Fictitious Cross-Play
GAIL Generative Adversarial Imitation Learning
GRF Google Research Football
HULHE Heads-Up Limit Texas Hold’em
HUNL Heads-Up No-Limit Texas Hold’em
JPSRO Joint Policy-Space Response Oracle
LCFR Linear Counterfactual Regret Minimization
LLM Large Language Model
MARL Multi-Agent Reinforcement Learning
MCCFR Monte Carlo Counterfactual Regret Minimization
MCTS Monte Carlo Tree Search
MDP Markov Decision Process
MG Markov Game
ML Machine Learning
MOBA Multiplayer Online Battle Arena
MSS Meta-Strategy Solver
NE Nash Equilibrium
NeuPL Neural Population Learning
NFSP Neural Fictitious Self-Play
ODO Online Double Oracle
PBR Preference-based Best Response
PBS Public Belief State
PFSP Prioritized Fictitious Self-Play
POMG Partially Observable Markov Game
PSRO Policy-Space Response Oracle
RCFR Regression Counterfactual Regret Minimization
ReBeL Recursive Belief-based Learning
RL Reinforcement Learning
R-NaD Regularized Nash Dynamics
RTS Real-Time Strategy
SD-CFR Single Deep Counterfactual Regret Minimization
SPAG Self-Play of Adversarial Game
SPIN Self-Play fIne-tuNing
TSP Traveling Salesman Problem
URR UnRestricted-Restricted
XDO Extensive-Form Double Oracle

### v4

@@ 5.5. Real-World Applications
Self-play is effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks (Lewis et al., 2017). Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) (Wang et al., 2024). Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse and Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
@@
Deploying self-play directly in the physical world remains challenging. Its iterative, trial‑and‑error nature is computationally expensive and can produce behaviors that are impractical or even unsafe outside a controlled environment. Consequently, most recent work performs self‑play in simulators, and practical adoption hinges on bridging the simulation‑to‑reality (Sim2Real) gap. When this gap is small, self‑play has already proved useful. For instance, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020). Moreover, EvoPlay (Wang et al., 2023) leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator (Jumper et al., 2021) to narrow the Sim2Real gap. In multi-robot systems, self-play has been recently utilized for competitive robot sports tasks like quadruped and humanoid football (Liu et al., 2022b; Haarnoja et al., 2024; Xiong et al., 2024; Su et al., 2025a), robot-arm and humanoid table tennis (DAmbrosio et al., 2025; Su et al., 2025b), and multi-drone volleyball (Xu et al., 2025b; Zhang et al., 2025), with substantial efforts dedicated to Sim2Real transitions for real-world success. A complementary research direction is to boost the sample efficiency of self-play so that policies can be trained online, reducing or even eliminating the need for elaborate Sim2Real pipelines in real-world deployments.
6. Conclusion
@@
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play, that agents iteratively refine their policies by interacting with historical or concurrent versions of themselves or other evolving agents, has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the game theory background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. We provide detailed explanations on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. We then compare the four categories within this common lens and draw design insights that can guide future self-play research. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within various scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Also, future work can focus on how to integrate self-play with LLMs and achieve real-world applications. In conclusion, self-play is vital to modern RL research, offering key insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.